OpenStudy (anonymous):

After 8 years, Ofelia earned $980 in simple interest from a CD into which she initially deposited $7000. What was the annual interest rate of the CD?

OpenStudy (anonymous):

The simple interest formula is;  So when we substitute in values that we know we get; \[980 = (7000) (r) (8)\] So now we need to find rate, we use this; \[r = \frac{ \left( 980 \right) }{ \left( 7000 \right)\left( 8 \right) }\] What is the value of r, or the rate?